OpenStudy (anonymous):

(8r - 3) - (7r + 1) = -6 Solve and check your solution. So far I have (8r - 7r) - (-3 + 1) = -6 r - 2 = -6 r = -4 Check (8[-4] - 3) - (7[-4] + 1) = -6 (-32 - 3) - (-28 + 1) = -6 -35 - (-27) does not = -6 HELP!!!!!

hero (hero):

(8r - 3) - (7r + 1) = -6 8r - 3 - 7r - 1 = -6 8r - 7r - 3 - 1 = -6 r - 4 = -6 r = -6 + 4 r = - 2

hero (hero):

Check: (8r - 3) - (7r + 1) = -6 (8(-2) - 3) -(7(-2)+1) = -6 (-16 - 3) -(-14 + 1) = -6 (-19)+ (14 - 1) = -6 -19 + 13 = -6 -6 = -6

OpenStudy (anonymous):

Reworking the problem as we speak, thanks for your help Hero your a big help.